OpenStudy (askme12345):

Can someone check my answer?! (How to solve is already answered) (:

OpenStudy (reemii):

a) sum the values, divide by 10. (empirical mean \(\frac{1}{n}\sum_{i=1}^{n} x_i\))

OpenStudy (askme12345):

doesnt that just give me the average @reemii

OpenStudy (reemii):

that's right. its' fine. a "point" estimate is a real number computed from the data. usually people choose the mean \(\bar x\).

OpenStudy (reemii):

(b) from http://en.wikipedia.org/wiki/Student's_t-distribution#Confidence_intervals This : \[\overline{X}_n - A\frac{S_n}{\sqrt{n}}, \overline{X}_n + A\frac{S_n}{\sqrt{n}}\] is the confidence interval, where \(\bar X_n\) is the mean computed in (a) \(S_n\) is the empirical standard deviation \(A\) is the 99.5% quantile of the student distribution.